OpenStudy (anonymous):

exact value for cos^-1(cos5pi/4)

OpenStudy (y2o2):

5pi/4

OpenStudy (nikvist):

\[\arccos{\left(\cos\frac{5\pi}{4}\right)}=\frac{3\pi}{4}\]